Check the convergence of the sequence defined by 𝑒𝑛+1 = 1 2 (𝑒𝑛 + π‘Ž 𝑒𝑛 ) , π‘Ž > 0. Note that this is the sequence associated with finding the square root of a number π‘Ž > 0 by the Newton’s method


1
Expert's answer
2021-11-03T08:49:48-0400

Let the sequenceΒ "u_n"Β converges toΒ "b."Β .

Then,


"u_{n+1}=\frac{1}{2}(u_n+\frac{a}{u_n})"

"b=\frac{1}{2}(b+\frac{a}{b})\\"

Solving this equation,we get


"b=\frac{1}{2}(b+\frac{a}{b})\\"

"2b^2=b^2+a""b^2=a"

"b=\pm\sqrt{a}"

Given that "a>0." Then "b>0=>b=\sqrt{a}."

Therefore the sequence defined by

"u_{n+1}=\frac{1}{2}(u_n+\frac{a}{u_n}), a>0"

converges to "\sqrt{a}."